# remarkRemark

### v1

@@ 3 Fast sinc quadrature for large time evaluation
We first consider (P1) without regard for (P2). A basic approach is to consider the worst-case scenario where $t=T$, and then discretize the integral in (10) with as many points as is needed. Due to the $e^{-i\omega t}$ term, the number of points required to achieve any accuracy must scale linearly with $T$. An alternative approach, which turns out to be equivalent to the “scaled convolution” method presented in [2], is readily suggested by Shannon–Whittaker–Kotelnikov Sampling Theorem [44, 45]. We present the idea from the perspective of Shannon sampling, as this usefully links the decay behavior of $u(x,t)$ to smoothness and analyticity properties of $\hat{U}$.
The sampling theorem states that in the Payey-Wiener space consisting of functions $f$ such that $f$ is continuous on $\mathbb{R}$ with bounded $L_{2}$ norm and
@@
We cannot sample $u(x,t)$ directly since it is what we are trying to compute. To convert (15) to a practical method, we note that if $\hat{U}(x,\omega)$ were truly bandlimited, then the Fourier coefficients of the periodic extension of $\hat{U}(x,\cdot)$ on $[W_{1},W_{2}]$ can exactly be expressed using equally-spaced samples of $u(x,t)$ over $\mathbb{R}$. Letting $y\in[0,1]$, we apply the change of variables $\omega=Py+W_{1}$, where $P=W_{2}-W_{1}$. There are coefficient functions $\{c_{-m}(x),\cdots,c_{m}(x)\}$ such that
@@
The coefficients can be computed for any fixed $x$ by sampling $\hat{U}(x,\cdot)$ on $12m\!+\!$ 1 equally-spaced points over $[W_{1},W_{2})$ and then applying the FFT. In the truly compactly supported case, it follows from an inspection of (6) that
@@
Once the coefficients are known, the fast sinc transform [22, 29] can be used to efficiently evaluate $u_{m}(x,t)$ at $N$ points $\{t_{1},\ldots,t_{N}\}\subset[0,T]$ at each fixed point $x_{k}\in X$. The computational cost is $\mathcal{O}(P\log P+m+N)$ operations, and the FINUFFT-based implementation [29] applies batching techniques that substantially improve the practical run time when (18) must be evaluated for many values of $x$.
The approximation in (18) is equivalently derived by convolving $u(x,t)$ with a scaled sinc function. This is perhaps the simplest windowing approximation strategy among many possibilities, including the use of Kaiser-Bessel functions, Gaussians, bump functions, or prolate-spheriodal wave functions. Alternative strategies could potentially improve the rate at which (18) converges to $u(x,t)$ as $m$ grows [25], but we leave this for future investigation. We emphasize that the primary advantage of (18) over basic quadrature choices is that the $t$-dependent oscillatory content in the integrand is reduced to a multiplicative factor (outside the integrand) that does not directly dictate the rate at which $\hat{U}(x,\cdot)$ must be sampled. This solves (P1).
@@ 3.1 Convergence of the truncated sinc expansion
We now turn to (P2). For the sake of analysis, we assume in this section that $\hat{U}(x,\cdot)$ is truly compactly supported on $[W_{1},W_{2}]$. The fundamental challenge in using (18) is that $\hat{U}(x,\cdot)$ becomes difficult to approximate with trigonometric polynomials if the decay of $|u(x,t)|$ is very slow. This in turn is closely connected to the smoothness and analyticity properties of $\hat{U}(x,\cdot)$. We succinctly state the obvious coupling between the decay rate of $|u(x,t)|$ and the convergence of $u_{m}\to u$ in the following lemma, which is a direct consequence of the Shannon–Whittaker–Kotelnikov Sampling Theorem and (17).
@@
For a fixed $\tilde{x}\in\Omega$, let $\hat{U}(\tilde{x},\omega)$ be a solution to (4) that is compactly supported on $[W_{1},W_{2}]$. Let $u(\tilde{x},t)=\mathcal{F}^{-1}(\hat{U}(\tilde{x},\omega))$ and define $u_{m}(\tilde{x},t)$ as in (18). Let $10<\epsilon<$ 1. Suppose there is a monotonically decreasing and invertible function $g(t)$ such that $|u(x,t)|=\Theta(g(t))$. Then,
@@
Lemma 3.1 implies that we can only guarantee that $\|u(\tilde{x},\cdot)-u_{m}(\tilde{x},\cdot)\|_{\infty,[0,\infty)}\leq\epsilon$ if $m$ is roughly on the order of a time point $t_{\epsilon}$, where $|u(x,t_{\epsilon})|<\epsilon$ and decays steadily thereafter for all $t>t_{\epsilon}$. For example, if $\mathbb{R}^{d}\setminus\Omega$ does not possess trapping regions (e.g., the scatterer forms a star domain [^3] A set $\mathcal{D}\subset\mathbb{R}^{d}$ is called a star domain if there exists $p_{0}\in\mathcal{D}$ such that for all $p\in\mathcal{D}$, the line segment joining $p_{0}$ and $p$ also lies in $\mathcal{D}$.) and $d$ is odd [^4] For even $d$ the situation is slightly more complicated [36]., then the celebrated result of Morawetz, Ralston and Strauss [37] proves that $g(t)$ is an exponential function. This implies that $u_{m}\to u$ uniformly and pointwise at an exponential rate. In such a setting, we expect the sinc-based method to work extremely well. However, we also expect that $T$ will never be required to be especially large in such a case.
@@
The claim made in [2] is that by using (18), the cost of evaluating the scattered field no longer appears to depend directly on the magnitude of the largest numerically relevant time $T$. However, Lemma 3.1 suggests that the choice of $m$ required to achieve a specified accuracy $\epsilon$ implicitly depends linearly on some $t_{\epsilon}\in[0,T]$. If we take $T$ to be $t_{\epsilon}$, then this appears to be a contradiction. However, a subtlety not to be overlooked is that if we choose $\epsilon$ such that $t_{\epsilon}<T$, we have $\mathcal{O}(\epsilon)$ error, even for $t\in[t_{\epsilon},T]$. This is a stark contrast with, say, Gauss-Legendre quadrature, where the error only decays meaningfully if the number of quadrature points used to discretize (6) is $\mathcal{O}(t)$, and thus in the worst case is $\mathcal{O}(T)$. It is in this sense that, as the authors of [2] state, the accuracy of $u_{m}(x,t)$ over $[0,T]$ is tunable.
What can we say about $g(t)$ in cases where there is trapping? The decay of $|u(x,t)|$ may be arbitrarily slow, and it cannot be succinctly described, though some asymptotics related to the poles of the scattering operator can be numerically extracted [12, 13]. We observe that the convergence of $u_{m}\to u$ in this setting can be so slow that (18) is no longer numerically useful.
@@ 4 A damping+correction scheme for strong trapping
As indicated in the last section, the decay rate of $|u(\tilde{x},t)|$ is an indicator of the convergence rate of $u\to u_{m}$. This is closely connected to the distance of poles of the meromorphic extension of $\hat{U}(\tilde{x},\cdot)$, which come closer to the real line as trapping grows more severe. A natural idea to improve matters is to consider a function with better analyticity properties by perturbing $\omega$ up into the complex plane. An early example of this applied in the context of scattering is found in [41]. Letting $\omega\in[W_{1},W_{2}]$, the function $\mathcal{F}^{-1}\hat{U}(x,\omega+\delta i)$ corresponds to a solution to the damped wave equation. Using contour integration and the fact that $\hat{U}(x,\cdot)$ is analytic in the upper half of the complex $\omega$-plane, we can use this function to recover the undamped solution without explicitly having to sample $\hat{U}(x,\cdot)$ on the real line.
We choose $\Gamma=\Gamma_{0}\cup\Gamma_{cR}\cup\Gamma_{\delta}\cup\Gamma_{cL}$ to form a rectangle. Specifically, $\Gamma_{0}$ is the line segment $[W_{1},W_{2}]$ traversed left to right, $\Gamma_{cR}$ is the line segment $[W_{2},W_{2}+\delta i]$ traversed bottom to top, $\Gamma_{\delta}$ is the line segment $[W_{1}+\delta i,W_{2}+\delta i]$ traversed left to right, and $\Gamma_{cL}$ is the line segment $[W_{1},W_{1}+\delta i]$ traversed top to bottom. Letting $I_{a}$ denote the contour integral over $\Gamma_{a}$ we conclude that (6) is equivalent to
@@ 4.1 Evaluating the damped term
Here, $\{\tilde{c}_{j}(x)\}_{j=1}^{m}$ are the positively-indexed Fourier coefficients of $\hat{U}(x,\omega+\delta i)$. These coefficients will decay more rapidly than those of $\hat{U}(\tilde{x},\cdot)$ so that $\tilde{u}_{m}\to u$ more rapidly than $u_{m}\to u$. While this improvement may seem modest when $\delta$ is small, it is enough to make a substantial difference in the number of required Helmholtz solves. Figure 4 illustrates this fact. The left picture shows profiles of the real part of solutions $\hat{U}(\tilde{x},\omega+i\delta)$ over $\omega\in[0.7,16]$ with $\tilde{x}=[0,0]^{T}$ selected as the center of the C-shaped keyhole scatterer shown in Figure 2. The right image shows the magnitude of the Fourier coefficients of $\hat{U}(\tilde{x},\omega+\delta i)$ for each $\delta$.
@@ 4.2 Evaluating the correction terms
[…]
@@ 5 Numerical Results
In this section we include several experiments that illustrate the capabilities of the solver. Videos of simulations can be found at the link listed in [46]. We have implemented the solver in MATLAB, with all code publicly available in [47]. We make use of several existing packages, including Flatiron’s FINUFFT library [8] and their implementation of the fast-sinc transform [29], the fmmlib2D [21] suite, and the Zetatrap package [48]. Specialized colormaps used in our visualizations are from the slanCM package [30]. Finally, we make use of the so-called chop function developed in Chebfun [3] for checking the resolution quality of (18) and enabling the adaptive selection of $m$. To improve the accessibility of our experiments, we include MATLAB code in Appendix A that generates all of the scatterers we use.
@@
Figure 6: The $2$-norm error of solutions (relative to a highly resolved solution) computed over three points in time and the 20 spatial points shown on the right (black dots) is plotted on a logarithmic scale against the parameter $\delta$. Each line shows the error behavior for a fixed choice of $m$ as $\delta$ grows.
@@ 5.1 Convergence behavior in a smooth domain
In this example, we examine the behavior of the solution as $m$ and $\delta$ are varied. The problem we consider involves a $C$-shaped scatterer with a smooth boundary. The incident wave is set as in (9) with $\omega_{0}=30$, $z_{0}=[1,0]^{T}$, and $t_{0}=20$. The frequency band over which $\hat{U}(x,\cdot)$ is relevant is $[21,38]$. Figure 5 shows the real part of the total at various points in time. Note that at $t=150$, the scattered field is still on the order of $10^{-2}$.
[…]
@@ 5.2 Multiply-connected domains
[…]
@@ 5.3 Domains with corners
Time-frequency-based methods possess two major potential advantages over methods that discretize directly in time. First, the former is not burdened by CFL-like restrictions that couple the spatial discretization to a fixed-size timestep. Second, the spatial discretization for solving a 2D problem only involves resolution along a 1D curve (the boundary of $\Omega$). The damping+correction fast-sinc method is effective for handling domains with both corners and trapping, which is a regime in which traditional time-stepping methods seriously struggle.
[…]
Figure 9: Magnitude of total field in a “fat radiator” domain at $t=31.311$ (left), $t=42.787$ (center) and $t=95.574$ (right)
@@ 5.3.1 Implementation details for nonsmooth domains
To handle corners, we apply a standard refinement scheme to the discretization of the integral equation (13) and the contour integral (11) that defines $\hat{U}(x,\omega)$. As discussed in [34], the scatterer is divided into several pieces, called panels, and Gauss-Legendre quadrature is applied to each panel with Kolm-Rokhlin correction [27] applied to manage singularities in the layer potentials. To improve accuracy, the panels are refined geometrically into the corners. This basic method is sufficient for general 2D domains and can also be adapted to smooth domains with regions of high curvature. Alternative methods include QBX-type quadratures [26] and other specialized corner quadrature schemes [11, 24, 43].

### v2

@@ 3 Fast sinc quadrature for large time evaluation
We first consider (P1) without regard for (P2). A basic approach is to consider the worst-case scenario where $t=T$, and then discretize the integral in (10) with as many points as is needed. Due to the $e^{-i\omega t}$ term, the number of points required to achieve any accuracy must scale linearly with $T$. An alternative approach, which turns out to be equivalent to the “scaled convolution” method presented in [anderson2020high], is readily suggested by Shannon–Whittaker–Kotelnikov Sampling Theorem [shannon1949communication, whittaker1915xviii]. We present the idea from the perspective of Shannon sampling, as this usefully links the decay behavior of $u(x,t)$ to smoothness and analyticity properties of $\hat{U}$.
The sampling theorem states that in the Paley-Wiener space consisting of functions $f$ such that $f$ is continuous on $\mathbb{R}$ with bounded $L_{2}$ norm and
@@
We cannot sample $u(x,t)$ directly since it is what we are trying to compute. To convert (15) to a practical method, we note that if $\hat{U}(x,\omega)$ were truly bandlimited, then the Fourier coefficients of the periodic extension of $\hat{U}(x,\cdot)$ on $W_{1},W_{2}]$ can exactly be expressed using equally-spaced samples of $u(x,t)$ over $\mathbb{R}$. Letting $y\in[0,1]$, we apply the change of variables $\omega=Py+W_{1}$, where $P=W_{2}-W_{1}$. There are coefficient functions $\{c_{-m}(x),\cdots,c_{m}(x)\}$ such that
@@
The coefficients can be computed for any fixed $x$ by sampling $\hat{U}(x,\cdot)$ on 2m+12m\!+\!1 equally-spaced points over $W_{1},W_{2})$ and then applying the FFT. In the truly compactly supported case, it follows from an inspection of (6) that
@@
Once the coefficients are known, the fast sinc transform [greengard2007fast, lawrenceFS] can be used to efficiently evaluate $u_{m}(x,t)$ at $N$ points $\{t_{1},\ldots,t_{N}\}\subset[0,T]$ at each fixed point $x_{k}\in X$. The computational cost is $\mathcal{O}(P\log P+m+N)$ operations, and the FINUFFT-based implementation [lawrenceFS] applies batching techniques that substantially improve the practical run time when (18) must be evaluated for many values of $x$.
The approximation in (18) is equivalently derived by convolving $u(x,t)$ with a scaled sinc function. This is perhaps the simplest windowing approximation strategy among many possibilities, including the use of Kaiser-Bessel functions, Gaussians, bump functions, or prolate-spheriodal wave functions. Alternative strategies could potentially improve the rate at which (18) converges to $u(x,t)$ as $m$ grows [kircheis2024fast], but we leave this for future investigation. We emphasize that the primary advantage of (18) over basic quadrature choices is that the $t$-dependent oscillatory content in the integrand is reduced to a multiplicative factor (outside the integrand) that does not directly dictate the rate at which $\hat{U}(x,\cdot)$ must be sampled. This solves (P1).
@@ 3.1 Convergence of the truncated sinc expansion
We now turn to (P2). For the sake of analysis, we assume in this section that $\hat{U}(x,\cdot)$ is truly compactly supported on $W_{1},W_{2}]$. The fundamental challenge in using (18) is that $\hat{U}(x,\cdot)$ becomes difficult to approximate with trigonometric polynomials if the decay of $|u(x,t)|$ is very slow. This in turn is closely connected to the smoothness and analyticity properties of $\hat{U}(x,\cdot)$. We succinctly state the obvious coupling between the decay rate of $|u(x,t)|$ and the convergence of $u_{m}\to u$ in the following lemma, which is a direct consequence of the Shannon–Whittaker–Kotelnikov Sampling Theorem and (17).
@@
For a fixed $\tilde{x}\in\Omega$, let $\hat{U}(\tilde{x},\omega)$ be a solution to (4) that is compactly supported on $W_{1},W_{2}]$. Let $u(\tilde{x},t)=\mathcal{F}^{-1}(\hat{U}(\tilde{x},\omega))$ and define $u_{m}(\tilde{x},t)$ as in (18). Let $0<\epsilon<1$. Suppose there is a monotonically decreasing and invertible function $g(t)$ such that $|u(x,t)|=\Theta(g(t))$. Then,
@@
Lemma 3.1 implies that we can only guarantee that $\|u(\tilde{x},\cdot)-u_{m}(\tilde{x},\cdot)\|_{\infty,[0,\infty)}\leq\epsilon$ if $m$ is roughly on the order of a time point $t_{\epsilon}$, where $|u(x,t_{\epsilon})|<\epsilon$ and decays steadily thereafter for all $t>t_{\epsilon}$. For example, if $\mathbb{R}^{d}\setminus\Omega$ does not possess trapping regions (e.g., the scatterer forms a star domain [^3] A set $\mathcal{D}\subset\mathbb{R}^{d}$ is called a star domain if there exists $p_{0}\in\mathcal{D}$ such that for all $p\in\mathcal{D}$, the line segment joining $p_{0}$ and $p$ also lies in $\mathcal{D}$.) and $d$ is odd [^4] For even $d$ the situation is slightly more complicated [morawetz1975decay]., then the celebrated result of Morawetz, Ralston and Strauss [morawetz1977decay] proves that $g(t)$ is an exponential function. This implies that $u_{m}\to u$ uniformly and pointwise at an exponential rate. In such a setting, we expect the sinc-based method to work extremely well. However, we also expect that $T$ will never be required to be especially large in such a case.
@@
The claim made in [anderson2020high] is that by using (18), the cost of evaluating the scattered field no longer appears to depend directly on the magnitude of the largest numerically relevant time $T$. However, Lemma 3.1 suggests that the choice of $m$ required to achieve a specified accuracy $\epsilon$ implicitly depends linearly on some $t_{\epsilon}\in[0,T]$. If we take $T$ to be $t_{\epsilon}$, then this appears to be a contradiction. However, a subtlety not to be overlooked is that if we choose $\epsilon$ such that $t_{\epsilon}<T$, we have $\mathcal{O}(\epsilon)$ error, even for $t\in[t_{\epsilon},T]$. This is a stark contrast with, say, Gauss-Legendre quadrature, where the error only decays meaningfully if the number of quadrature points used to discretize (6) is $\mathcal{O}(t)$, and thus in the worst case is $\mathcal{O}(T)$. It is in this sense that, as the authors of [anderson2020high] state, the accuracy of $u_{m}(x,t)$ over $[0,T]$ is tunable.
What can we say about $g(t)$ in cases where there is trapping? The decay of $|u(x,t)|$ may be arbitrarily slow, and it cannot be succinctly described, though some asymptotics related to the poles of the scattering operator can be numerically extracted [bruno2025efficient, bruno2024evaluation]. We observe that the convergence of $u_{m}\to u$ in this setting can be so slow that (18) is no longer numerically useful.
@@ 4 A damping+correction scheme for strong trapping
[…]
We choose $\Gamma=\Gamma_{0}\cup\Gamma_{cR}\cup\Gamma_{\delta}\cup\Gamma_{cL}$ to form a rectangle. Specifically, $\Gamma_{0}$ is the line segment $W_{1},W_{2}]$ traversed left to right, $\Gamma_{cR}$ is the line segment $W_{2},W_{2}+\delta i]$ traversed bottom to top, $\Gamma_{\delta}$ is the line segment $W_{1}+\delta i,W_{2}+\delta i]$ traversed left to right, and $\Gamma_{cL}$ is the line segment $W_{1},W_{1}+\delta i]$ traversed top to bottom. Letting $I_{a}$ denote the contour integral over $\Gamma_{a}$ we conclude that (6) is equivalent to
@@ 4.1 Evaluating the damped term
Figure 4: Left: Plots of $\hat{U}(\tilde{x},\omega+\delta i)$ for $\delta=0$ (blue), $\delta\approx.005$ (red), $\delta\approx.015$ (yellow), and $\delta\approx.02$ (purple) over $\omega\in[.7,16]$, with $\tilde{x}$ selected as the center of the $C$-shaped curve in Figure 2. The plots are stacked vertically for illustrative purposes. Right: The magnitudes of the normalized Fourier coefficients for $\hat{U}(\tilde{x},\omega+\delta_{j}i)$ for each $\delta_{j}$. Faster decay implies that fewer terms are needed to compute the sum in (19) and thereby evaluate $I_{\delta}$.
@@ 4.2 Evaluating the correction terms
The damped terms are not oscillatory integrals and can be evaluated using Gauss-Legendre quadrature. They become increasingly easier to approximate as $\delta$ grows larger, so a natural impulse is to push $\delta$ to be as large as possible. However, exponential growth in the correction terms will cause numerical overflow if $\delta$ is too large, imposing a $T$-dependent limit on how far up $\delta$ can go. A heuristic for estimating this limit is that $\delta$ should not exceed $L(T)=\tau_{mach}\ln(2)/(150T)$, where where $2^{\tau_{mach}}$ is the largest representable floating point number for a given machine. For IEEE 754 double-precision, $\tau_{mach}=1024$, and $L(T)\approx 6.826\ln(2)/T$. As an example, for $T=200$, $L(T)\approx.0237$. For $T=500$, $L(T)\approx.0095$.
@@ 5 Numerical Results
In this section we include several experiments that illustrate the capabilities of the solver. Videos of simulations can be found at the link listed in [HOMEyoutube]. We have implemented the solver in MATLAB, with all code publicly available in [TFAW]. We make use of several existing packages, including Flatiron’s FINUFFT library [barnett2018non] and their implementation of the fast-sinc transform [lawrenceFS], the fmmlib2D [greengard2020fmmlib2d] suite, and the Zetatrap package [wu2021zeta]. Specialized colormaps used in our visualizations are from the slanCM package [Slanpack]. Finally, we make use of the so-called chop function developed in Chebfun [aurentz2017chopping] for checking the resolution quality of (18) and enabling the adaptive selection of $m$. To improve the accessibility of our experiments, we include MATLAB code in Appendix A that generates all of the scatterers we use.
@@
Figure 6: The 22-norm error of solutions (relative to a highly resolved solution) computed over three points in time and the 20 spatial points shown on the right (black dots) is plotted on a logarithmic scale against the parameter $\delta$. Each line shows the error behavior for a fixed choice of $m$ as $\delta$ grows.
@@ 5.1 Convergence behavior in a smooth domain
In this example, we examine the behavior of the solution as $m$ and $\delta$ are varied. The problem we consider involves a $C$-shaped scatterer with a smooth boundary. The incident wave is set as in (9) with $\omega_{0}=30$, $z_{0}=[1,0]^{T}$, and $t_{0}=20$. The frequency band over which $\hat{U}(x,\cdot)$ is relevant is $[21,38]$. Figure 5 shows the geometry and the real part of the total field at various points in time. Note that at $t=150$, the scattered field is still on the order of $10^{-2}$.
The images in Figure 5 were produced with a highly resolved approximation to the solution with $m=4000$ and $\delta=.02$. In Figure 6, we look at the error between this solution and solutions computed with different choices of $\delta$ and $m$. We measure the 2-norm error calculated over $20$ points in the domain (see Figure 6) and five equally-spaced points in time taken over the interval $[5,150]$.
@@ 5.2 Multiply-connected domains
No special modifications are required to handle multiply-connected domains. In Figure 7 we show a solution involving two offset crescents that “pass” the wavefield between one-another. For this example the incident wave is chosen as in (9) with $\omega_{0}=10$, $z_{0}=[0,1]^{T}$, and $t_{0}=20$. The frequencies over which the problem is relevant are $\omega\in[7,12]$, and we set $\delta=.02$ with $m=850$.
@@ 5.3 Domains with corners
Time-frequency-based methods possess two major potential advantages over methods that discretize directly in time. First, the former is not burdened by CFL-like restrictions that can provide highly restrictive bounds on the maximal time-step allowed, in particular in situations that involve mesh refinement. Second, the spatial discretization for solving a 2D problem only involves resolution along a 1D curve (the boundary of $\Omega$). The damping+correction fast-sinc method is effective for handling domains with both corners and trapping, which is a regime in which traditional time-stepping methods seriously struggle.
We consider two domains: the absolute value of the total field is shown in a keyhole domain in Figure 8, and in a more challenging magnetron domain in Figure 9. A different magnetron domain with thinner channels is shown in Figure 1. For these experiments, $\delta=.025$ and $m=1000$.
Figure 9: Magnitude of total field in a magnetron domain at $t=31.311$ (left), $t=42.787$ (center) and $t=95.574$ (right)
@@ 5.3.1 Implementation details for nonsmooth domains
To handle corners, we apply a standard refinement scheme to the discretization of the integral equation (13) and the contour integral (11) that defines $\hat{U}(x,\omega)$. As discussed in [martinsson2019fast], the scatterer is divided into several pieces, called panels, and Gauss-Legendre quadrature is applied to each panel with Kolm-Rokhlin correction [kolm2001numerical] applied to manage singularities in the layer potentials. To improve accuracy, the panels are refined geometrically into the corners. This basic method is sufficient for general 2D domains and can also be adapted to smooth domains with regions of high curvature. Alternative methods include QBX-type quadratures [klockner2013quadrature] and other specialized corner quadrature schemes [bruno2009high, hoskins2019numerical, serkh2019solution].
5.4 A 3D example
Since the damping+correction scheme only involves contour integration with respect to the frequency parameter $\omega$, it can be applied straightforwardly when $\Omega$ is a 3D domain. Nothing about the inverse Fourier transform changes, but it is substantially more computationally intensive to evaluate $\hat{U}(x,\omega)$. As in the 2D setting, boundary integral methods can be used to solve for each density function associated with a given frequency, so that with the FMM, one can quickly evaluate $\hat{U}(x_{k},\omega_{j})$. We use the FMMLIB3D package [greengard2020fmmlib3d]. Specialized fast quadrature schemes for surface integrals are needed to work with such domains. Our example in Figure 10 is a proof-of-concept that uses a biperiodic “cruller” domain from the BIE3D package [BIE3d] and Zeta-correction quadrature [wu2021corrected]. In future work, we plan to implement 3D examples involving more sophisticated domains with severe trapping.
Figure 10: A biperiodic 3D “cruller” scatterer (left), and the real part of the total field at times $t=18.923$ (center) and $t=23.156$ (right).